OpenStudy (anonymous):

The sum of two numbers is 51. Their difference is 13. What is the smaller of the two numbers? A. 18 B. 19 C. 20 D. 21

OpenStudy (anonymous):

do you want to learn the material or just get the answer?

OpenStudy (anonymous):

that's not correct

hero (hero):

x + y = 51 x - y = 13 2x = 64 x = 64/2 x = 32 y = 51- 32 y = 19 19 is the smaller number.